OpenStudy (anonymous):

What is the GCF of 45, 60, and 330? PLZ help me!

Mehek (mehek14):

Do you know what GCF is?

Mehek (mehek14):

@aahmed123 ?

OpenStudy (anonymous):

:P i can help

Mehek (mehek14):

First, start by listing the factors of each number

OpenStudy (anonymous):

Greatest Common Factor

Mehek (mehek14):

\(\Large{45:}\) \(\Large{60:}\) \(\Large{330:}\)

Mehek (mehek14):

start listing the factors

OpenStudy (anonymous):

i mite as well give you it straight

OpenStudy (anonymous):

ok i list all the factors down

Mehek (mehek14):

yes

Mehek (mehek14):

@firesonic101 you can't do that. It's against the rules

OpenStudy (anonymous):

i know

Mehek (mehek14):

then don't

OpenStudy (anonymous):

ok i wont

Mehek (mehek14):

@aahmed123 I'll help you start with 45

Mehek (mehek14):

so what times 45 equals 45?

OpenStudy (anonymous):

2,025

OpenStudy (anonymous):

o.o what

Mehek (mehek14):

no 25 * 2 = 50....

Mehek (mehek14):

1, 45 is the first one

OpenStudy (anonymous):

wait sorry :I

OpenStudy (anonymous):

ugh my heads going bonkers REALLY sorry

Mehek (mehek14):

it's ok you will get a decimal so skip on to 3

Mehek (mehek14):

45/3?

OpenStudy (anonymous):

45/3=15

Mehek (mehek14):

ok you have 1, 3, 15, 45 now try 4 if it gives you a decimal, skip on to 5

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

45/5=9

Mehek (mehek14):

1, 3, 5, 9, 15, 45 that's all of them

Mehek (mehek14):

now move on to 60 start by 1 and go on

OpenStudy (anonymous):

ok so i should do the same with all the numbers??

Mehek (mehek14):

yes

OpenStudy (anonymous):

ok give me a sec

OpenStudy (anonymous):

OK im back!

Mehek (mehek14):

so what did you get?

OpenStudy (anonymous):

well i know what GCF means and i found that 15 was the GCF, am i correct?

Mehek (mehek14):

yes that's correct

OpenStudy (anonymous):

thank you!

Mehek (mehek14):

yw :)